Calculate the knee angle in the following participant performing deep water running given the following coordinates: Hip (10,10), Knee (3, 6), Ankle (8, 4).
Answer #1

hip vector=10i+10j

Knee vector=3i+6j

Ankle vector=8i+4j

a=the vector between the knee and hip=10i+10j-(3i+6j)=7i+4j

b=the vector between the knee and ankle=10i+10j-(8i+4j)=2i+6j

angle b/w vector a and b=